# Computing the Elementary Symmetric Polynomials in Positive Characteristics

### v2

@@ 2 Order-$2$ zero space of the elementary symmetric polynomials
Hence, we conclude that $e_{d-1}^{n}(a)=0$. We further notice that, for every $i\in[n]$,
$e_{d-1}^{n}(a)=\frac{\partial e_{d}^{n}}{\partial x_{i}}(a)+a_{i}\frac{\partial e_{d-1}^{n}}{\partial x_{i}}(a),$
so we conclude that either $a_{i}=0$ or $\frac{\partial e_{d-1}^{n}}{\partial x_{i}}(a)=0$. Because, for every $i\in I$ we know that $a_{i}\neq 0$, we conclude that $\frac{\partial e_{d-1}^{n}}{\partial x_{i}}(a)=0$. Thus, we conclude that $(a_{i}\mid i\in I)\in V_{2}(e_{d-1}^{|I|})$. Hence, if we let $\nu_{I}:\mathbb{F}^{|I|}\rightarrow\mathbb{F}^{n}$ naturally add zeros in the indices not in $I$, we conclude that $V_{2}(e_{d}^{n})\cap\mathcal{S}_{I}\subseteq\nu_{I}(V_{2}(e_{d-1}^{|I|}))$. Noticing that $\nu_{I}$ does not change the dimension of the underlying variety,
$\dim(\mathcal{S}_{I}\cap V_{2}(e_{d}^{n}))\leq\dim(\nu_{I}(V_{2}(e_{d-1}^{|I|})))=\dim V_{2}(e_{d-1}^{|I|})\leq d-2,$
by Lemma 2.4. This completes the proof. ∎
@@ 3 The Symmetric Model
In this section, we will study a computational model defined from the elementary symmetric polynomials. Specifically, we will define Sym to represent the set of homogeneous polynomials of some degree, say $d$, that can be written as $e_{d}^{m}(L_{1},\ldots,L_{m})$, for some linear (homogeneous) polynomials $L_{1},\ldots,L_{m}$. We will then say that $\Sigma^{[k]}\text{Sym}$ are all polynomials that can be written as a linear combination of $k$ elements of Sym. We will use $\overline{\text{Sym}}$ and $\overline{\Sigma^{[k]}\text{Sym}}$ to denote the border versions of these classes. Later in the section, we will prove Theorem 3.9, from which we conclude that the fact that some polynomials cannot be computed by $\overline{\Sigma^{[k]}\text{Sym}}$ implies that they cannot be computed by $\overline{\Sigma^{[k]}\Pi\Sigma}$ circuits.
This model of computation was introduced and studied in [Shp02], where it was shown that, in fields of characteristic zero, every polynomial can be computed by Sym. This result used the fact that, under such conditions, every polynomial has finite Waring rank (the smallest $k$ such that a polynomial can be written as $L_{1}^{d}+\dots+L_{k}^{d}$, for linear $L_{i}$). This was proved in Lemma 2.4 of [Shp02], but we provide here a slight variation of this lemma, which is slightly stronger and uses a slightly different method.
@@
Unfortunately, the Waring rank model is known to be not universal in positive characteristic. If $\operatorname{char}(\mathbb{F})=p\neq 0$, then we observe that, for every $d\geq p$, monomials in $L^{d}$, where $L\in\mathbb{F}[x_{1},\ldots,x_{n}]$ is linear, must be divisible by some $x_{i}^{p}$. For example, we cannot represent multilinear polynomials from this model. In this section, we will build on this fact to show that $\overline{\text{Sym}}$ and $\overline{\Sigma^{[k]}\text{Sym}}$ are not universal. Specifically, we will show that polynomials in these classes can be expressed as the sum of a small number of reducible polynomials and an arbitrary number of powers of linear forms. Then, if we consider the computation of a multilinear polynomial, we can ignore these linear powers and only consider the reducible polynomials.
We will spend this section proving Theorem 1.2. Then, Theorem 1.3 will follow from Theorem 3.9. To begin with, we will restrict our attention to fields of characteristic two, as this will make the proofs simpler. We will focus on how to extend them to higher field characteristics after this.
@@ 3.1 The case of characteristic two
First, suppose that $f$ is in Sym. Let $L_{1},\ldots,L_{m}\in\mathbb{F}[x_{1},\ldots,x_{n}]$ be linear such that $e_{3}^{m}(L_{1},\ldots,L_{m})=f$. Then, by the Newton identities (8) and the fact that $p_{2}=e_{1}^{2}$, we observe that
@@
Now, let $g$ be a reducible homogeneous degree three polynomial and $q_{1},\ldots,q_{k}$ be linear. Let $g_{1}$ be homogeneous degree two and $g_{2}$ be linear such that $g=g_{1}g_{2}$. By Claim 3.2, let $L_{1},\ldots,L_{m}\in\mathbb{F}[x_{1},\ldots,x_{n}]$ be linear such that $g_{1}=e_{2}^{m}(L_{1},\ldots,L_{m})$ and $0=e_{1}^{m}(L_{1},\ldots,L_{m})$. Then, observe that
@@
As a simple corollary to this claim, we can use this to characterize elements of $\Sigma^{[k]}\text{Sym}$.
@@
Utilizing this characterization, we will now focus on determining a counter-example. We will focus our attention on multilinear polynomials, as, as we have previously described, they cannot be computed by cubes of linear forms in characteristic two. In fact, cubes of linear forms cannot contain any multilinear monomials, so we can ignore this part of the representation. Hence, it will suffice to find a polynomial that cannot be written as the multilinear part of a reducible polynomial. Specifically, we will focus on the polynomial $x_{1}x_{2}x_{3}+x_{4}x_{5}x_{6}$.
The argument will follow by showing that, if the multilinear part of a reducible polynomial has nonzero coefficients of $x_{1}x_{2}x_{3}$ and $x_{4}x_{5}x_{6}$, there must be another multilinear monomial with a nonzero coefficient. We will do this by constructing a polynomial in the coefficients of a linear and quadratic polynomial that relates the coefficients of the multilinear monomials in their product. Then, the fact that this also works in the border case will follow from the polynomial identity.
@@
The polynomial $f=x_{1}x_{2}x_{3}+x_{4}x_{5}x_{6}\in\mathbb{F}[x_{1},\ldots,x_{6}]$ is not in Sym. This is also true in the border setting.
@@
Suppose that $f$ is in Sym. Let $g$ be a reducible homogeneous degree three polynomial and $q_{1},\ldots,q_{k}$ be linear such that $f=g+q_{1}^{3}+\dots+q_{k}^{3}$. First, we observe that all monomials in $q_{1}^{3}+\dots+q_{k}^{3}$ are not multilinear. Thus, the coefficients of the multilinear monomials in $f$ are given by $g$.
@@
Our goal will be to use the fact that $c_{123}\neq 0$ and $c_{456}\neq 0$ (meaning that $c_{123}c_{456}\neq 0$) to show that some other $c_{ijk}\neq 0$, a contradiction. We will consider these $c_{ijk}$ values as polynomials in the coefficients of $g_{1}$ and $g_{2}$. Notice that we can conclude that there is a term within $c_{123}c_{456}$ that is non-zero, say $a_{1}b_{23}a_{4}b_{56}\neq 0$. Observe that this is generated by exactly one other partition of variables, namely $c_{234}c_{156}$, implying that this term is canceled when we take $c_{123}c_{456}+c_{234}c_{156}$. We can then apply the process repeatedly until we hit a point where the equation computes zero. We then observe that this is a metapolynomial in the coefficients of $f$, so this also applies in the border setting.
We will state this more formally to show that this indeed happens. Let $\mathcal{P}_{3}[6]$ represent all of the ways to partition $[6]$ into two sets of size three. Then, we will show that
$\sum_{\mathcal{I}\in\mathcal{P}_{3}[6]}\sum_{\{i,j,k\}\in\mathcal{I}}c_{ijk}\equiv 0\mod 2.$ (4)
@@
$\sum_{\mathcal{I}\in\mathcal{P}_{3}[6]}\sum_{\{i,j,k\}\in\mathcal{I}}c_{ijk}=2\sum_{\{i,i^{\prime}\}\subseteq[6]}\sum_{\{\{j,k\},\{j^{\prime},k^{\prime}\}\}\in\mathcal{P}_{2}([6]\setminus\{i,i^{\prime}\})}a_{i}a_{i^{\prime}}b_{jk}b_{j^{\prime}k^{\prime}}.\qed$
With a bit of carefulness, we can extend this argument to $\Sigma^{[k]}\text{Sym}$. We will use a similar polynomial to our counter-example, but we will, this time, add more monomials of the form $x_{i}y_{i}z_{i}$. We will then use the same identity as the previous case, but we will instead take a sum over partitions of the $3\ell$ variables into groups of three. We can now consider this as a polynomial in the coefficients in the decompositions of the reducible polynomials. Now, each monomial can be split up into $\ell$ parts based on the reducible polynomial the variable comes from. Then, we can permute the corresponding linear parts, as we did in the previous part, for those that came from the same equation, and we conclude that the number of partitions that generate this monomial is given by the product of factorials of the corresponding size of the groups. Thus, if $\ell$ is big enough, one of these sizes must be at least two, and we conclude that the whole equation is zero.
@@ 3.2 Higher characteristics
$=\sum_{i=1}^{p-1}(-1)^{i-1}e_{k-i}^{m}(L_{1},\ldots,L_{m})p_{i}^{m}(L_{1},\ldots,L_{m})+(e_{1}^{m}(L_{1},\ldots,L_{m}))^{p+1}+L_{1}^{p+1}+\dots+L_{m}^{p+1}\qed$
Notice that this claim is not as strong as the claim for characteristic two. Currently, it is not known how to extend it, but only one directions is necessary for the rest of the proofs. The biggest part of the problem is that it is not clear how to extend Claim 3.2 to this case, even if the field characteristic is three.
@@
Consider a fixed monomial $x_{i_{1}}\dots x_{i_{p+1}}$, and let $M=\{i_{1},\ldots,i_{p+1}\}$. Observe that the coefficient of this monomial is not influenced by the term $q_{1}^{p+1}+\dots+q_{N}^{p+1}$. We will abuse notation and sometimes write, given a subset $S\subseteq[n]$, $a_{S}^{(i)}$ or $b_{S}^{(i)}$ to represent the coefficient corresponding to the multilinear monomial given by the set $S$. Thus, we have that the coefficient of this monomial in $f$ is given by
$\sum_{i=1}^{k}\sum_{S\in{M\choose d_{i}}}a_{S}^{(i)}b_{M\setminus S}^{(i)}=c_{M}.$
@@
$F=\sum_{\mathcal{I}\in\mathcal{P}_{p+1}[n]}\prod_{I\in\mathcal{I}}c_{I}\equiv 0\mod p.$
Once we show this, the proof is clearly complete.
We will fix a monomial in $F$. Observe that it can be written as $a_{S_{1}}^{(j_{1})}b_{S_{1}^{\prime}}^{(j_{1})}\dots a_{S_{\ell}}^{(j_{\ell})}b_{S_{\ell}^{\prime}}^{(j_{\ell})}$, where $S_{1}\cup S_{1}^{\prime}\cup\dots\cup S_{\ell}\cup S_{\ell}^{\prime}=[n]$ and $t_{1},\ldots,t_{\ell}\in[k]$. Notice that the coefficient of this monomial is equal to the number of choices of $\mathcal{I}$ that could generate it. In fact, a choice of $\mathcal{I}$ is a valid selection if and only if there is a bijection $\pi:[\ell]\rightarrow[\ell]$ such that $\mathcal{I}=\{S_{1}\cup S_{\pi(1)}^{\prime},\ldots,S_{\ell}\cup S_{\pi(\ell)}^{\prime}\}$ and $t_{i}=t_{\pi(i)}$. For $i\in[k]$, let $n_{i}$ be the number of $t_{j}$ values such that $t_{j}=i$. Notice that $n_{1}+\dots+n_{k}=\ell$. Then, the number of such selections of bijections is exactly equal to $n_{1}!\cdots n_{k}!$. Because $\ell>k\cdot(p-1)$, there is a $i\in[k]$ such that $n_{i}\geq p$. Thus, we have shown that the coefficient of this monomial is zero modulus $p$. This then completes the proof by the same argument as the proof of Claim 3.5. ∎

### v3

@@ 2 Order-$2$ zero space of the elementary symmetric polynomials
Hence, we conclude that $e_{d-1}^{n}(a)=0$. We further notice that, by (7), for every $i\in[n]$,
$e_{d-1}^{n}(a)=a_{i}\frac{\partial e_{d-1}^{n}}{\partial x_{i}}(a).$
Thus, we observe that $f_{d}(a)=e_{d-1}^{n}(a)=0$ and, for $i\leq n-k$, we have that $\frac{\partial f_{d}}{\partial x_{i}}(a)=\frac{\partial e_{d-1}^{n}}{\partial x_{i}}(a)=0$. Hence, we conclude, by definition, that $(a_{1},\ldots,a_{n-k})\in V_{2}(e_{d-1}^{n-k})$.
Interestingly, we can also get a full understanding when the degree is large in comparison to the number of variables.
Claim 2.9.
If n<2d1n<2d-1, then we have that $\dim V_{2}(e_{d}^{n})=d-2$.
Proof.
Recall that $\mathcal{P}_{k}([n])$ denotes the set of all partitions of $[n]$ into $k$ groups. We thus know that $V_{2}(e_{d}^{n})\subseteq\bigcup_{\pi\in\mathcal{P}_{d-1}([n])}P_{\pi}$ from the proof of Lemma 2.4 (see (4)). Let $V\subseteq V_{2}(e_{d}^{n})$ be an irreducible component (as $V_{2}(e_{d}^{n})$ is a variety), and let $\pi$ be such that $V\subseteq P_{\pi}$, which we know exists as $P_{\pi}$ is a vector space, meaning it is irreducible. Further notice that $\dim P_{\pi}=d-1$, so, to show that $\dim V\leq d-2$, it suffices to show that $V\subsetneq P_{\pi}$.
Let $\pi=\{S_{1},\ldots,S_{k}\}$ and $L_{i}=|S_{i}|$ for every $i\in[k]$. Without loss of generality, we assume that $i\in S_{i}$. Observe that $P_{\pi}\cong\mathbb{F}^{k}$ via the isomorphism $\varphi:P_{\pi}\rightarrow\mathbb{F}^{k}$ defined by $\varphi(x_{1},\ldots,x_{n})=(x_{1},\ldots,x_{k})$. After applying the morphism to $V_{2}(e_{d}^{n})$, we see that, in this space, it is defined by the equations
$e_{d}^{n}(P_{\pi})=\sum_{\overset{(j_{1},\ldots,j_{k})\in[L_{1}]\times\dots\times[L_{k}]}{j_{1}+\dots+j_{k}=d}}{L_{1}\choose j_{1}}\dots L_{k}\choose j_{k}x_{1}^{j_{1}}\dots x_{k}^{j_{k}},$
and, for $i\in[k]$,
$\frac{\partial e_{d}^{n}}{\partial x_{i}}(P_{\pi})=\sum_{\overset{(j_{1},\ldots,j_{k})\in[L_{1}]\times\dots\times[L_{i}-1]\times\dots\times[L_{k}]}{j_{1}+\dots+j_{k}=d-1}}{L_{1}\choose j_{1}}\dots L_{i}-1\choose j_{i}\dots L_{k}\choose j_{k}x_{1}^{j_{1}}\dots x_{k}^{j_{k}}.$
To separate $V$ and $P_{\pi}$, it suffices to show that one of these polynomials is not identically zero. Our strategy will be to find $\ell_{1},\ldots,\ell_{c}\in[k]$ such that $L_{\ell_{1}}+\dots+L_{\ell_{c}}\in\{d-1,d\}$. Without loss of generality, we say that $\ell_{i}=i$ for every $i\in[c]$. Notice that, if we can find such sets, then the coefficient of $x_{1}^{L_{1}}\dots x_{c}^{L_{c}}$ in at least one of these polynomials is one.
It now suffices to show that such a selection exists. We will describe a simple algorithmic selection procedure. The idea is to greedily grab the biggest set iteratively that does not cause the sum of the set sizes we have taken to be larger than $d$. We then must only argue that this procedure will always end with the sum being d1d-1 or $d$. To see this, notice that, because $d$ is close to $n$, there should be a lot of sets of size one. Hence, from this strategy, the proof follows from the following claim.
Claim 2.10.
Let $N$ be the number of sets of size exactly one. Then, for each $S_{i}$, we have that $N\geq L_{i}-2$.
Proof.
By assumption, we know that $n\leq 2d-2$. Further, because all of the other sets are of size at least two, we conclude that $N+2(d-1-N-1)+L_{i}\leq n$. Thus, it follows that $2d-2\geq 2d-N+L_{i}-4$. Hence, we have that $N\geq L_{i}-2$. ∎
Thus, we can guarantee that there is enough $1$’s to get to d1d-1. ∎
Observe that this does not characterize when the answer is d1d-1 versus d2d-2. Specifically, we do not know what happens when $p\neq 0$, $n\geq 2d-1$, and $n-d+1\equiv 0\mod p$, but $n-d+1\not\equiv 0\mod p^{\lceil\log_{p}{d}\rceil}$. Based on the behavior in smaller cases, we propose the following characterization.
Conjecture 2.11.
Given $1\leq d\leq n$ and $\operatorname{char}(\mathbb{F})=p$, we have that $\dim V_{2}(e_{d}^{n})=d-1$ if and only if
  • $p\neq 0$,
  • $n\geq 2d-1$, and
  • $n-d+1\equiv 0\mod p^{\lceil\log_{p}{d}\rceil}$.
Otherwise, $\dim V_{2}(e_{d}^{n})=d-2$.
@@ 3 The Symmetric Model
In this section, we will study a computational model defined from the elementary symmetric polynomials. Specifically, we will define Sym to represent the set of homogeneous polynomials of some degree, say $d$, that can be written as $e_{d}^{m}(L_{1},\ldots,L_{m})$ for some linear (homogeneous) polynomials $L_{1},\ldots,L_{m}$. We will then say that $\Sigma^{[k]}\text{Sym}$ are all polynomials that can be written as a linear combination of $k$ elements of Sym. We will use $\overline{\text{Sym}}$ and $\overline{\Sigma^{[k]}\text{Sym}}$ to denote the border versions of these classes. Later in the section, we will prove Theorem 3.9, meaning the fact that some polynomials cannot be computed by $\overline{\Sigma^{[k]}\text{Sym}}$ implies that they cannot be computed by $\overline{\Sigma^{[k]}\Pi\Sigma}$ circuits.
This model of computation was introduced and studied in [29], where it was shown that, in fields of characteristic zero, every polynomial can be computed by Sym. This result used the fact that, under such conditions, every polynomial has finite Waring rank (the smallest $k$ such that a polynomial can be written as $L_{1}^{d}+\dots+L_{k}^{d}$, for linear $L_{i}$). This was proved in Lemma 2.4 of [29], but we provide here a slight variation of this lemma, which is slightly stronger and uses a slightly different method.
@@
Unfortunately, the Waring rank model is known to be not universal in positive characteristic. If $\operatorname{char}(\mathbb{F})=p\neq 0$, then we observe that, for every $d\geq p$, monomials in $L^{d}$, where $L\in\mathbb{F}[x_{1},\ldots,x_{n}]$ is linear, must be divisible by some $x_{i}^{p}$. For example, we cannot represent multilinear polynomials using this model. In this section, we will build on this fact to show that $\overline{\text{Sym}}$ and $\overline{\Sigma^{[k]}\text{Sym}}$ are not universal. Specifically, we will show that polynomials in these classes can be expressed as the sum of a small number of reducible polynomials and an arbitrary number of powers of linear forms. Then, if we consider the computation of a multilinear polynomial, we can ignore these linear powers and only consider the reducible polynomials.
We will spend this section proving Theorem 1.2. Then, Theorem 1.3 will follow from Theorem 3.9. To begin with, we will restrict our attention to fields of characteristic two, as this will make the proofs simpler. We will then extend them to higher field characteristics.
@@ 3.1 The case of characteristic two
First, suppose that $f$ is in Sym. Let $L_{1},\ldots,L_{m}\in\mathbb{F}[x_{1},\ldots,x_{n}]$ be linear such that $e_{3}^{m}(L_{1},\ldots,L_{m})=f$. Then, by the Newton identities (9) and the fact that $p_{2}^{m}=(e_{1}^{m})^{2}$, we observe that
@@
For the other direction, let $g$ be a reducible homogeneous degree three polynomial and $q_{1},\ldots,q_{k}$ be linear. Let $g_{1}$ be homogeneous degree two and $g_{2}$ be linear such that $g=g_{1}g_{2}$. By Claim 3.2, let $L_{1},\ldots,L_{m}\in\mathbb{F}[x_{1},\ldots,x_{n}]$ be linear such that $g_{1}=e_{2}^{m}(L_{1},\ldots,L_{m})$ and $0=e_{1}^{m}(L_{1},\ldots,L_{m})$. Then, observe that
@@
As a simple corollary to this claim, we can characterize the elements of $\Sigma^{[k]}\text{Sym}$.
@@
Utilizing this characterization, we will now focus on constructing a polynomial that cannot be computed under this model of computation. We will focus our attention on multilinear polynomials, as, as we have previously described, they cannot be computed by cubes of linear forms in characteristic two. In fact, cubes of linear forms cannot contain any multilinear monomials, so we can ignore this part of the representation. Hence, it will suffice to find a polynomial that cannot be written as the multilinear part of a reducible polynomial. Specifically, we will focus on the polynomial $x_{1}y_{1}z_{1}+x_{2}y_{2}z_{2}$.
The argument will follow by showing that, if the multilinear part of a reducible polynomial has nonzero coefficients of $x_{1}y_{1}z_{1}$ and $x_{2}y_{2}z_{2}$, there must be another multilinear monomial with a nonzero coefficient. We will do this by constructing a ’metapolynomial’ in the coefficients of a linear and quadratic polynomial that relates the coefficients of the multilinear monomials in their product. Then, the fact that this also works in the border case will follow from the polynomial identity.
@@
The polynomial $f=x_{1}y_{1}z_{1}+x_{2}y_{2}z_{2}\in\mathbb{F}[x_{1},\ldots,x_{6}]$ is not in Sym. This is also true in the border setting.
@@
For simplicity, we will write $f=x_{1}x_{2}x_{3}+x_{4}x_{5}x_{6}$. Suppose that $f$ is in Sym. Let $g$ be a reducible homogeneous degree three polynomial and $q_{1},\ldots,q_{k}$ be linear such that $f=g+q_{1}^{3}+\dots+q_{k}^{3}$. First, we observe that all monomials in $q_{1}^{3}+\dots+q_{k}^{3}$ are not multilinear. Thus, the coefficients of the multilinear monomials in $f$ are given by $g$.
@@
Our goal will be to use the fact that $c_{123}\neq 0$ and $c_{456}\neq 0$ (meaning that $c_{123}c_{456}\neq 0$) to show that some other $c_{ijk}\neq 0$, a contradiction. We will consider these $c_{ijk}$ values as polynomials in the coefficients of $g_{1}$ and $g_{2}$.
Notice that that there must be a coefficient (in terms of the $a$’s and $b$’s) within $c_{123}c_{456}$ that is non-zero, say $a_{1}b_{23}a_{4}b_{56}$, for example. This monomial also has a nonzero coefficient in $c_{234}c_{156}$ (which corresponds to a different variable partition). In fact, if we consider all possible partitioning of the variables, this is the only other partition that has this coefficient nonzero. Our strategy will then be to take combinations of these pairs formed from partitions, which will form a ’metapolynomial’ in the coefficient of $f$ that vanishes, meaning that the result will also work in the border setting.
We will state this more formally to show that this indeed happens. Let $\mathcal{P}_{3}[6]$ represent all of the ways to partition $[6]$ into two sets of size three. [^5] Note that the notation for partitions is slightly different from the notation used in previous sections. Then, we will show that
$\sum_{\mathcal{I}\in\mathcal{P}_{3}[6]}\prod_{\{i,j,k\}\in\mathcal{I}}c_{ijk}\equiv 0\mod 2.$ (5)
@@
$\sum_{\mathcal{I}\in\mathcal{P}_{3}[6]}\prod_{\{i,j,k\}\in\mathcal{I}}c_{ijk}=2\sum_{\{i,i^{\prime}\}\subseteq[6]}\sum_{\{\{j,k\},\{j^{\prime},k^{\prime}\}\}\in\mathcal{P}_{2}([6]\setminus\{i,i^{\prime}\})}a_{i}a_{i^{\prime}}b_{jk}b_{j^{\prime}k^{\prime}}.\qed$
With a bit of carefulness, we can extend this argument to $\Sigma^{[k]}\text{Sym}$. We will use a similar polynomial for our counterexample, but we will add more monomials of the form $x_{i}y_{i}z_{i}$. We will then use the same identity as the previous case, but we will instead take a sum over partitions of the $3\ell$ variables into groups of three. We can now consider this as a polynomial in the coefficients of the decompositions of the reducible polynomials. Now, each monomial can be split up into $\ell$ parts based on the reducible polynomial the variable comes from. Then, we can permute the corresponding linear parts, as we did in the previous part, for those that came from the same equation, and we conclude that the number of partitions that generate this monomial is given by the product of factorials of the corresponding size of the groups. Thus, if $\ell$ is big enough, one of these sizes must be at least two, and we conclude that the whole equation is zero.
@@ 3.2 Higher characteristics
$=\sum_{i=2}^{p}(-1)^{i-1}e_{p+1-i}^{m}(L_{1},\ldots,L_{m})p_{i}^{m}(L_{1},\ldots,L_{m})+(e_{1}^{m}(L_{1},\ldots,L_{m}))^{p+1}+L_{1}^{p+1}+\dots+L_{m}^{p+1}\qed$
Notice that this claim is not as strong as the claim for characteristic two. Currently, it is not known how to extend it, but only one directions is necessary for the rest of the proofs. The biggest part of the problem is that it is not clear how to extend Claim 3.2 to this case, even if the field characteristic is just three.
@@
Consider a fixed monomial $x_{i_{1}}\dots x_{i_{p+1}}$, and let $M=\{i_{1},\ldots,i_{p+1}\}$. Observe that the coefficient of this monomial is not influenced by the term $q_{1}^{p+1}+\dots+q_{N}^{p+1}$, so we will ignore these terms. We will abuse notation and sometimes write, given a subset $S\subseteq[n]$, $a_{S}^{(i)}$ or $b_{S}^{(i)}$ to represent the coefficient corresponding to the multilinear monomial given by the set $S$. Thus, we have that the coefficient of this monomial in $f$ is given by
$\sum_{i=1}^{k}\sum_{S\in{M\choose d_{i}}}a_{S}^{(i)}b_{M\setminus S}^{(i)},$
which we will denote by $c_{M}$.
@@
$F=\sum_{\mathcal{I}\in\mathcal{P}_{p+1}[n]}\prod_{I\in\mathcal{I}}c_{I}\equiv 0\mod p,$
where $\mathcal{P}_{p+1}[n]$ denotes the set of all ways to partition $[n]$ into groups of size $p+1$. Once we show this, the proof is clearly complete.
We will fix a monomial in $F$. Observe that it can be written as $a_{S_{1}}^{(j_{1})}b_{S_{1}^{\prime}}^{(j_{1})}\dots a_{S_{\ell}}^{(j_{\ell})}b_{S_{\ell}^{\prime}}^{(j_{\ell})}$, where $S_{1}\cup S_{1}^{\prime}\cup\dots\cup S_{\ell}\cup S_{\ell}^{\prime}=[n]$ and $t_{1},\ldots,t_{\ell}\in[k]$. Notice that the coefficient of this monomial is equal to the number of choices of $\mathcal{I}\in\mathcal{P}_{p+1}[n]$ that could generate it. In fact, a choice of $\mathcal{I}$ is a valid selection if and only if there is a bijection $\pi:[\ell]\rightarrow[\ell]$ such that each $S\in\mathcal{I}$ can be written as $S_{i}\cup S_{\pi(i)}^{\prime}$ for some $i\in\ell$ and we have that $t_{j}=t_{\pi(j)}$ for all $j\in[\ell]$. For $i\in[k]$, let $n_{i}$ be the number of $t_{j}$ values such that $t_{j}=i$. Notice that $n_{1}+\dots+n_{k}=\ell$. Then, the number of such selections of bijections is exactly equal to $n_{1}!\cdots n_{k}!$. Because $\ell>k\cdot(p-1)$, there is a $i\in[k]$ such that $n_{i}\geq p$. Thus, we have shown that the coefficient of this monomial is zero modulus $p$. This then completes the proof by the same argument as the proof of Claim 3.5. ∎